OpenStudy (anonymous):

what is 2 times the square root of 3 over three and all squared

OpenStudy (debbieg):

\[\Large \left( \dfrac{ 2\sqrt{3} }{ 3} \right)^2\] Use: \(\Large \left( \dfrac{ a }{ b} \right)^2=\dfrac{ a^2 }{ b^2}\) and also: \(\Large (ab)^m=a^m\cdot b^m\) You'll have some reducing to do at the end.

OpenStudy (debbieg):

And welcome to Open Study! :)